# A Bag of Tricks for Few-Shot Class-Incremental Learning

### v1

@@ Abstract
We present a bag of tricks framework for few-shot class-incremental learning (FSCIL), which is a challenging form of continual learning that involves continuous adaptation to new tasks with limited samples. FSCIL requires both stability and adaptability, i.e., preserving proficiency in previously learned tasks while learning new ones. Our proposed bag of tricks brings together eight key and highly influential techniques that improve stability, adaptability, and overall performance under a unified framework for FSCIL. We organize these tricks into three categories: stability tricks, adaptability tricks, and training tricks. Stability tricks aim to mitigate the forgetting of previously learned classes by enhancing the separation between the embeddings of learned classes and minimizing interference when learning new ones. On the other hand, adaptability tricks focus on the effective learning of new classes. Finally, training tricks improve the overall performance without compromising stability or adaptability. We perform extensive experiments on three benchmark datasets, CIFAR-100, CUB-200, and miniIMageNet, to evaluate the impact of our proposed framework. Our detailed analysis shows that our approach substantially improves both stability and adaptability, establishing a new state-of-the-art by outperforming prior works in the area. We believe our method provides a go-to solution and establishes a robust baseline for future research in this area.
Keywords:
Few-shot Learning Continual Learning Class-incremental Learning
@@ 1 Introduction
Continual learning is a machine learning paradigm that focuses on the ability of a model to learn new knowledge, without forgetting what it previously learned. In real-world applications, machine learning models often encounter scenarios where they must adapt to novel classes with only a limited number of samples available for learning. This scenario has inspired the introduction of an exciting paradigm called Few-Shot Class Incremental Learning (FSCIL) [34]. Existing literature [34, 40] has demonstrated that traditional continual learning approaches are ineffective in FSCIL, primarily due to the scarcity of labelled data during incremental learning sessions. This data limitation often leads to overfitting on the novel classes, resulting in a well-known catastrophic forgetting issue.
@@
Some prior works have linked the problem of catastrophic forgetting with high adaptability (or plasticity) during incremental training [8, 38]. Consequently, these approaches reduce the adaptability by utilizing an incremental-frozen framework, where the encoder is trained only in the base session and remains frozen during the incremental sessions [38, 34, 8, 6]. The resulting methods provide very high stability, but very little adaptability. This phenomenon is often referred to as the stability-adaptability dilemma [24, 39] in the FSCIL, where high stability causes reduced adaptability and vice-versa.
@@
In this work, we combine a collection of techniques under a bag of tricks framework with the goal of concurrently enhancing both the stability and adaptability of FSCIL. These tricks have never been explored together and their impact on stability, adaptability, and overall performance has not been studied under the same framework. We categorize the techniques into three main groups: (i) stability, (ii) adaptability, and (iii) training. Stability tricks improve the separation among ‘base’ classes within the learned embedding space, which minimizes interference with previously learned classes when introducing novel classes during the incremental session. This results in greater stability and reduced forgetting of already learned classes. These approaches include incorporating supervised contrastive loss [15], pre-assigning prototypes [36], and including pseudo-classes [17] during training. On the other hand, adaptability tricks enhance the model’s ability to learn novel classes in the incremental session with careful tuning of the encoder through incremental fine-tuning [33] and SubNet tuning [14]. Finally, training tricks include training a larger encoder, including a pre-training step [4], and adding a pre-text task [11] in order to help boost the overall performance without compromising either stability or adaptability.
@@
Extensive experiments on three popular benchmark datasets, namely CIFAR100 [16], miniImageNet [30] and CUB200 [35], demonstrate the effectiveness of our framework. Overall, the proposed framework improves the performance by 3.42%, 1.25%, and 2.46% on CIFAR-100, CUB-200, and miniImageNet, respectively. Incorporating these tricks not only surpasses the overall accuracy of existing FSCIL methods but also demonstrates improved stability (reduced forgetting), higher adaptability (improvements on novel classes), and better separation in the embedding space. Study on the representations learned under stability tricks demonstrates lower intra-class variance and higher inter-class distance between the learned classes, which ensures lower forgetting and thus higher training stability and overall accuracy. Further study on adaptability tricks shows substantial improvement in the performance on new classes while retaining performance for the learned ones. We also present an ablation study to understand the impact of stability, adaptability, and training tricks on overall performance. This study highlights that stability tricks have the most significant impact, followed by adaptability tricks. Since one of our training tricks involves utilizing a larger encoder, we report the performance of previous state-of-the-art methods on the larger encoder and also report the performance of our method on the smaller encoder to ensure fair comparisons. Following prior works, we report the main results with 5-shot setting experiments on commonly used benchmark datasets. Moreover, to study the behaviour of our method under data-scarce scenarios, we also experiment with 1-shot and 2-shot settings. Finally, to evaluate scalability with respect to a number of classes, we also report our results on the large-scale ImageNet-1K [29] dataset with 1000 classes.
Overall, we make the following contributions:
@@ 2.1 Class Incremental Learning
Class-incremental learning (CIL) is a continual learning paradigm that involves continuously learning novel classes while preserving the knowledge of previously learned ones [21]. Existing literature on CIL can be broadly categorized into three main groups. The first group, commonly referred to as the replay-based method, stores past samples in a memory bank for rehearsal during incremental sessions to ensure retention of old knowledge [26, 27]. The second group is regularization-based, focusing on preventing significant changes in the parameters to prevent forgetting [18, 20]. In contrast, the third group dynamically expands the network architecture to accommodate the learning of novel classes [42, 32]. In practice, the model is trained over several sessions, with each session introducing novel classes to learn. In traditional CIL, each session includes an adequate number of labelled samples for every novel class, as well as the option to store some of the previous samples for rehearsal in future incremental sessions. However, as per the definition and problem setup of FSCIL, neither of these assumptions holds.
@@ 2.2 Few-Shot Class-Incremental Learning
In real-world applications, assuming that an incremental session contains a large number of samples for each novel class is impractical [34]. FSCIL [34, 12, 7, 19] tackles this challenging scenario where the model needs to incrementally learn novel classes with only a few samples per class. It also assumes that no samples from previous sessions are available, which causes privacy concerns in many domains. Consequently, none of the traditional CIL methods, mentioned in the previous section, perform well in the FSCIL setting [40, 34]. However, there have been developments in FSCIL literature, which can be discussed in two broad categories. The first category trains the model only on the base session, keeping the model frozen in the incremental session [43, 31, 37]; while the second group tunes the model in the incremental sessions  [34, 6, 9, 38].
The main idea of the first category (frozen encoder-based methods) is to ensure greater separability of base classes in the embedding space so that novel classes in the incremental step can easily fit in this space with minimal interference [43, 31, 37]. This is often referred to as forward compatibility [33]. Prior works in the literature proposed different techniques aiming to enable forward compatibility. For instance, FACT [40] proposed to use virtual prototypes to force the embedding of different classes to be maximally separated while respecting the relative positioning of classes in the embedding space. SAVC [33] created virtual classes during the base session, and [36] assigned random prototypes that are maximally separated from one another. These methods offer relatively high stability, but little adaptability for incremental training.
The second group of methods tunes the encoder in the incremental session to provide better flexibility for learning new tasks, thus providing better adaptability. For example, MgSvF [38] strategically updated different components at different rates, effectively balancing the adaptation to new knowledge and the preservation of old ones. The exemplar relation distillation framework [9] constructed and updated an exemplar relation graph to facilitate the learning of novel classes. SoftNet [14] proposed to utilize the concept of the lottery-ticket hypothesis to find a sub-network of important parameters from the previous session, which is left frozen during incremental tuning with the rest of the parameters. A common problem of this second group of methods is that adaptability comes at the cost of stability. That is, the model’s performance in the old classes deteriorates as it learns novel classes. Overall, there is a lack of balance between the stability and adaptability in existing FSCIL methods, which we aim to improve with our bag of tricks.
@@ 3 Method
3.1 Preliminaries
In FSCIL, a model is trained across $T$ consecutive sessions, with each session introducing novel classes for the model to learn. Training data for each session $t\in T$ is labelled, $\mathcal{D}_{train}^{t}=\{(x_{i},~{}y_{i})\}_{i=0}^{N_{t}}$, where $x_{i}$ and $y_{i}$ are the $i$-th sample and the corresponding label. In FSCIL, only the base session (first session) contains a sufficient amount of samples for effective training. Subsequent incremental sessions contain only a few samples per class, typically organized in an $N$-way $K$-shot format, containing $K$ training samples for each of the $N$ classes. By definition, each session exclusively contains samples from novel classes, meaning that the label space for each session ($\mathcal{C}^{t}$) is mutually exclusive with others. The performance of a method is evaluated after each session on the test set $\mathcal{D}_{test}^{t}$, which contains samples from all classes encountered so far.
@@ 3.2 Baseline
We consider the incremental frozen framework as our baseline due to its proven effectiveness [37, 31, 33] in addressing the data-scarce incremental learning scenario of FSCIL. In this framework, the model $\phi(x)$ is trained only on the base session ($\mathcal{D}_{train}^{0}$) using a standard cross-entropy loss. The model $\phi(x)$ consists of an encoder, $f_{\theta}(x)\in\mathbb{R}^{d\times 1}$ and a classifier head $W\in\mathbb{R}^{d\times|\mathcal{C}^{0}|}$. The prediction can therefore be expressed as $\phi(x)=W^{T}f_{\theta}(x)$. After training on the base session, the encoder $f(x)$ remains frozen during the incremental sessions. To classify novel classes, the classifier $W$ is expanded with the classifier weight for novel classes parameterized by the prototype of each class. A prototype is the average of the embeddings of all samples belonging to that class, ${w}_{c}^{t}=\frac{1}{n_{c}^{t}}{\sum_{i=1}^{n_{c}^{t}}{f_{\theta}({{x}}_{c,i})}}$.
@@ 3.3 Stability Tricks
Stability tricks in our framework revolve around the idea that better separation of ‘base’ classes ensures improved stability in learning new classes in the incremental sessions [33]. As illustrated in Figure 1, better separation of base classes in the embedding space allows the novel classes to be placed in the embedding space without interfering with existing classes. This involves increasing the distance between classes (inter-class distance) while reducing the distance between samples within the same class (intra-class distance). This approach is also known as forward compatibility in FSCIL literature. Accordingly, we incorporate three techniques that can effectively promote better stability: training with a supervised contrastive loss, pre-assigning prototypes, and including pseudo-classes.
@@
Supervised Contrastive Loss. While most of the existing literature on FSCIL use the standard cross-entropy loss for learning during the base session, some prior works [33] have demonstrated that cross-entropy does not effectively separate classes in the embedding space. Some studies have indicated that the supervised contrastive loss (SupCon) [15] exhibits better separability in the embedding space. SupCon is a variation of the popular contrastive loss [5], which additionally includes class labels to guide representation learning in a supervised manner. Specifically, SupCon learns representations by pulling the samples (and their augmentations) of the same class closer in the embedding space while pushing the samples of different classes apart, resulting in a more separable embedding space than cross-entropy. In other words, SupCon forces the representation of each sample to be close to its corresponding class prototype (center of embedding), while pushing the prototypes away from one another. For a batch of labelled samples $\{(x_{i},~{}y_{i})\}_{i=0}^{N}$, the SupCon loss can be represented as:
[…]
@@
where $z=f_{\theta}(x)$, $N$ is the batch-size, and $N{y_{i}}$ is the number of positive samples from class $y_{i}$. The indicator function denoted by $\mathbf{1}{y_{i}=y_{j}}$ yields a value of 1 when indices $i$ and $j$ correspond to instances of the same class.
@@
Pre-assigning Prototypes. In SupCon, prototypes are learned along with the optimization of the model. However, it was shown in [36] that ‘pre-assigning’ prototypes that are maximally separated from one another ensures improved separation in the embedding space. In [36], the maximally separated prototypes were defined as simplex Equiangular Tight Frame (ETF) [23], a geometric arrangement of $K$ vectors in $d$-dimensional Euclidean space [36]. Each vector in this space has an Euclidean norm of 1, with any pair of distinct vectors yielding an inner product of $-1/(K-1)$. This specific inner product value corresponds to the largest possible angle between any two vectors in this space. Our implementation of the prototype pre-assignment slightly differs from [36] in two ways. First, we do not rely solely on the cross-entropy loss but instead incorporate the SupCon loss introduced earlier. Second, we do not assign the ETF prototypes randomly at the very beginning of training. Instead, we train the encoder with the SupCon for a few epochs and take advantage of the prototypes naturally formed by SupCon. Then, for each class, we assign the ETF prototype that is most closely aligned with its naturally formed prototype.
@@
Including Pseudo-classes. Previous studies such as FACT [40] and SAVC [33] introduced the concept of integrating pseudo-classes during the base session to serve as placeholders in the embedding space for novel classes. In SAVC, the pseudo-classes were generated by a pre-defined transformation, which was considered a more fine-grained variant of the original classes. In our work, we take a similar approach of including pseudo-classes. More specifically, we incorporate the pseudo-class generation concept of [17], which employs hard augmentations to transform the semantics of the sample and consider it as a pseudo-class.

### v2

@@ Abstract
We present a bag of tricks framework for few-shot class-incremental learning (FSCIL), which is a challenging form of continual learning that involves continuous adaptation to new tasks with limited samples. FSCIL requires both stability and adaptability, i.e., preserving proficiency in previously learned tasks while learning new ones. Our proposed bag of tricks brings together six key and highly influential techniques that improve stability, adaptability, and overall performance under a unified framework for FSCIL. We organize these tricks into three categories: stability tricks, adaptability tricks, and training tricks. Stability tricks aim to mitigate the forgetting of previously learned classes by enhancing the separation between the embeddings of learned classes and minimizing interference when learning new ones. On the other hand, adaptability tricks focus on the effective learning of new classes. Finally, training tricks improve the overall performance without compromising stability or adaptability. We perform extensive experiments on three benchmark datasets, CIFAR-100, CUB-200, and miniIMageNet, to evaluate the impact of our proposed framework. Our detailed analysis shows that our approach substantially improves both stability and adaptability, establishing a new state-of-the-art by outperforming prior works in the area. We believe our method provides a go-to solution and establishes a robust baseline for future research in this area.
@@ 1 Introduction
Continual learning is a machine learning paradigm that focuses on the ability of a model to learn new knowledge, without forgetting what it previously learned. In real-world applications, machine learning models often encounter scenarios where they must adapt to novel classes with only a limited number of samples available for learning. This scenario has inspired the introduction of an exciting paradigm called Few-Shot Class Incremental Learning (FSCIL) (Tao et al., 2020). Existing literature (Tao et al., 2020; Zhou et al., 2022a) has demonstrated that traditional continual learning approaches are ineffective in FSCIL, primarily due to the scarcity of labelled data during incremental learning sessions. This data limitation often leads to overfitting on the novel classes, resulting in a well-known catastrophic forgetting issue.
@@
Some prior works have linked the problem of catastrophic forgetting with high adaptability (or plasticity) during incremental training (Chi et al., 2022; Zhao et al., 2021). Consequently, these approaches reduce the adaptability by utilizing an incremental-frozen framework, where the encoder is trained only in the base session and remains frozen during the incremental sessions (Zhao et al., 2021; Tao et al., 2020; Chi et al., 2022; Cheraghian et al., 2021a). The resulting methods provide very high stability, but very little adaptability. This phenomenon is often referred to as the stability-adaptability dilemma (Peng et al., 2022b; Zhao et al., 2023) in the FSCIL, where high stability causes reduced adaptability and vice-versa.
@@
In this work, we combine a collection of techniques under a bag of tricks framework with the goal of concurrently enhancing both the stability and adaptability of FSCIL. These tricks have never been explored together, and their impact on stability, adaptability, and overall performance has not been studied under the same framework. We categorize the techniques into three main groups: (i) stability, (ii) adaptability, and (iii) training. Stability tricks improve the separation among ‘base’ classes within the learned embedding space, which minimizes interference with previously learned classes when introducing novel classes during the incremental session. This results in greater stability and reduced forgetting of already learned classes. These approaches include incorporating supervised contrastive loss (Khosla et al., 2020), pre-assigning prototypes (Yang et al., 2022), and including pseudo-classes (Lee et al., 2020) during training. On the other hand, adaptability tricks enhance the model’s ability to learn novel classes in the incremental session with careful tuning of the encoder through incremental SubNet tuning (Kang et al., 2023). Finally, training tricks include a pre-training step (Chen et al., 2020), and adding a pre-text task (Gidaris et al., 2018) in order to help boost the overall performance without compromising either stability or adaptability. In summary, the key innovation of our framework is combining a set of modules (tricks), which improve stability, adaptability, and overall performance at the same time. While we did not propose any new components, no prior work has explored such modules (tricks) together in the context of FSCIL under a unified framework.
@@
Extensive experiments on three popular benchmark datasets, namely CIFAR100 (Krizhevsky et al., 2009), miniImageNet (Russakovsky et al., 2015) and CUB200 (Wah et al., 2011), demonstrate the effectiveness of our framework. Overall, the proposed framework improves the performance by 3.22%, 1.1%, and 2.0% on CIFAR-100, CUB-200, and miniImageNet, respectively. Incorporating these tricks not only surpasses the overall accuracy of existing FSCIL methods but also demonstrates improved stability (reduced forgetting), higher adaptability (improvements on novel classes), and better separation in the embedding space. Study on the representations learned under stability tricks demonstrates lower intra-class variance and higher inter-class distance between the learned classes, which ensures lower forgetting and thus higher training stability and overall accuracy. Further study on adaptability tricks shows substantial improvement in the performance on new classes while retaining performance for the learned ones. We also present an ablation study to understand the impact of stability, adaptability, and training tricks on overall performance. This study highlights that stability tricks have the most significant impact, followed by adaptability tricks. Following prior works, we report the main results with 5-shot setting experiments on commonly used benchmark datasets. Moreover, to study the behaviour of our method under data-scarce scenarios, we also experiment with 1-shot and 2-shot settings. Additionally, our framework shows similar improvement for larger encoders like ViT, and larger ResNets. Finally, to evaluate scalability with respect to a number of classes, we also report our results on the large-scale ImageNet-1K (Russakovsky et al., 2009) dataset with 1000 classes. Overall, we make the following contributions:
@@ 2.1 Class Incremental Learning
Class-incremental learning (CIL) is a continual learning paradigm that involves continuously learning novel classes while preserving the knowledge of previously learned ones (Masana et al., 2022). Existing literature on CIL can be broadly categorized into three main groups. The first group, commonly referred to as the replay-based method, stores past samples in a memory bank for rehearsal during incremental sessions to ensure retention of old knowledge (Rebuffi et al., 2017; Rolnick et al., 2019). The second group is regularization-based, focusing on preventing significant changes in the parameters to prevent forgetting (Li & Hoiem, 2017; Liu et al., 2018). In contrast, the third group dynamically expands the network architecture to accommodate the learning of novel classes (Zhu et al., 2021a; Shi et al., 2022). In practice, the model is trained over several sessions, with each session introducing novel classes to learn. In traditional CIL, each session includes an adequate number of labelled samples for every novel class, as well as the option to store some of the previous samples for rehearsal in future incremental sessions. However, as per the definition and problem setup of FSCIL, neither of these assumptions holds.
@@ 2.2 Few-Shot Class-Incremental Learning
In real-world applications, assuming that an incremental session contains a large number of samples for each novel class is impractical (Tao et al., 2020). FSCIL (Tao et al., 2020; Hersche et al., 2022; Cheraghian et al., 2021b; Liu et al., 2022) tackles this challenging scenario where the model needs to incrementally learn novel classes with only a few samples per class. It also assumes that no samples from previous sessions are available, which causes privacy concerns in many domains. Consequently, none of the traditional CIL methods, mentioned in the previous section, perform well in the FSCIL setting (Zhou et al., 2022a; Tao et al., 2020). However, there have been developments in FSCIL literature, which can be discussed in two broad categories. The first category trains the model only on the base session, keeping the model frozen in the incremental session (Zhu et al., 2021b; Shi et al., 2021; Zhang et al., 2021); while the second group tunes the model in the incremental sessions  (Tao et al., 2020; Cheraghian et al., 2021a; Dong et al., 2021; Zhao et al., 2021).
The main idea of the first category (frozen encoder-based methods) is to ensure greater separability of base classes in the embedding space so that novel classes in the incremental step can easily fit in this space with minimal interference (Zhu et al., 2021b; Shi et al., 2021; Zhang et al., 2021). This is often referred to as forward compatibility (Song et al., 2023). Prior works in the literature proposed different techniques aiming to enable forward compatibility. For instance, FACT (Zhou et al., 2022a) proposed to use virtual prototypes to force the embedding of different classes to be maximally separated while respecting the relative positioning of classes in the embedding space. SAVC (Song et al., 2023) created virtual classes during the base session, and (Yang et al., 2022) assigned random prototypes that are maximally separated from one another. These methods offer relatively high stability, but little adaptability for incremental training.
The second group of methods tunes the encoder in the incremental session to provide better flexibility for learning new tasks, thus providing better adaptability. For example, MgSvF (Zhao et al., 2021) strategically updated different components at different rates, effectively balancing the adaptation to new knowledge and the preservation of old ones. The exemplar relation distillation framework (Dong et al., 2021) constructed and updated an exemplar relation graph to facilitate the learning of novel classes. SoftNet (Kang et al., 2023) proposed to utilize the concept of the lottery-ticket hypothesis to find a sub-network of important parameters from the previous session, which is left frozen during incremental tuning with the rest of the parameters. A common problem of this second group of methods is that adaptability comes at the cost of stability. That is, the model’s performance in the old classes deteriorates as it learns novel classes. Overall, there is a lack of balance between the stability and adaptability in existing FSCIL methods, which we aim to improve with our bag of tricks.
@@ 3 Method
3.1 Overview
In FSCIL, a model is trained across $T$ consecutive sessions, with each session introducing novel classes for the model to learn. Training data for each session $t\in T$ is labelled, $\mathcal{D}_{train}^{t}=\{(x_{i},~{}y_{i})\}_{i=0}^{N_{t}}$, where $x_{i}$ and $y_{i}$ are the $i$-th sample and the corresponding label. In FSCIL, only the base session (first session) contains a sufficient amount of samples for effective training. Subsequent incremental sessions contain only a few samples per class, typically organized in an $N$-way $K$-shot format, containing $K$ training samples for each of the $N$ classes. By definition, each session exclusively contains samples from novel classes, meaning that the label space for each session ($\mathcal{C}^{t}$) is mutually exclusive with others. The performance of a method is evaluated after each session on the test set $\mathcal{D}_{test}^{t}$, which contains samples from all classes encountered so far. In our framework, we include a pre-training stage before the base training. Overall, the framework is divided into three training stages: pre-training, base training, and incremental training.
@@ 3.2 Baseline
We consider the incremental frozen framework as our baseline due to its proven effectiveness (Zhang et al., 2021; Shi et al., 2021; Song et al., 2023) in addressing the data-scarce incremental learning scenario of FSCIL. In this framework, the model $\phi(x)$ is trained only on the base session ($\mathcal{D}_{train}^{0}$) using a standard cross-entropy loss, $L_{ce}=(\phi(x),~{}y)$. The model $\phi(x)$ consists of an encoder, $f_{\theta}(x)\in\mathbb{R}^{d\times 1}$ and a classifier head $W\in\mathbb{R}^{d\times|\mathcal{C}^{0}|}$. The prediction can therefore be expressed as $\phi(x)=W^{T}f_{\theta}(x)$. After training on the base session, the encoder $f(x)$ remains frozen during the incremental sessions. To classify novel classes, the classifier $W$ is expanded with the classifier weight for novel classes parameterized by the prototype of each class. A prototype is the average of the embeddings of all samples belonging to that class, ${w}_{c}^{t}=\frac{1}{n_{c}^{t}}{\sum_{i=1}^{n_{c}^{t}}{f_{\theta}({{x}}_{c,i})}}$.
@@ 3.3 Stability Tricks
Stability tricks in our framework revolve around the idea that better separation of ‘base’ classes ensures improved stability in learning new classes in the incremental sessions (Song et al., 2023). As illustrated in Figure 1, better separation of base classes in the embedding space allows the novel classes to be placed in the embedding space without interfering with existing classes. This involves increasing the distance between classes (inter-class distance) while reducing the distance between samples within the same class (intra-class distance). This approach is also known as forward compatibility in FSCIL literature. Accordingly, we incorporate three techniques that can effectively promote better stability: training with a supervised contrastive loss, pre-assigning prototypes, and including pseudo-classes.
@@
Supervised Contrastive Loss. While most of the existing literature on FSCIL use the standard cross-entropy loss for learning during the base session, some prior works (Song et al., 2023) have demonstrated that cross-entropy does not effectively separate classes in the embedding space. Some studies have indicated that the supervised contrastive loss (SupCon) (Khosla et al., 2020) exhibits better separability in the embedding space. SupCon is a variation of the popular contrastive loss (Chen et al., 2020), which additionally includes class labels to guide representation learning in a supervised manner. Specifically, SupCon learns representations by pulling the samples (and their augmentations) of the same class closer in the embedding space while pushing the samples of different classes apart, resulting in a more separable embedding space than cross-entropy. In other words, SupCon forces the representation of each sample to be close to its corresponding class prototype (center of embedding), while pushing the prototypes away from one another. For a batch of labelled samples $\{(x_{i},~{}y_{i})\}_{i=0}^{N}$, the SupCon loss can be represented as:
$\begin{split}\mathcal{L}_{\text{{sup}}}=\sum_{i=1}^{N}\bigg{(}\frac{-1}{N_{y_{i}}-1}\sum_{j=1}^{N}\mathbf{1}_{[i\neq j]}\cdot\mathbf{1}_{[y_{i}=y_{j}]}\cdot log\frac{exp(z_{i}\cdot z_{j}/\tau)}{\sum_{k=1}^{N}\mathbf{1}_{[k\neq i]}exp(z_{i}\cdot z_{k}/\tau)}\bigg{)}~{},\end{split}$ (1)
@@
where $z=f_{\theta}(x)$, $N$ is the batch-size, and $N{y_{i}}$ is the number of positive samples from class $y_{i}$. The indicator function denoted by $\mathbf{1}_{[y_{i}=y_{j}]}$ yields a value of 1 when indices $i$ and $j$ correspond to instances of the same class, $\tau$ is a temperature parameter.
@@
Pre-assigning Prototypes. In SupCon, prototypes are learned along with the optimization of the model. However, it was shown in Yang et al. (2022) that ‘pre-assigning’ prototypes that are maximally separated from one another ensures improved separation in the embedding space. In Yang et al. (2022), the maximally separated prototypes were defined as simplex Equiangular Tight Frame (ETF) (Papyan et al., 2020) $P$, a geometric arrangement of $K$ vectors in $d$-dimensional Euclidean space (Yang et al., 2022). Each vector in this space has an Euclidean norm of 1, with any pair of distinct vectors yielding an inner product of $-1/(K-1)$. This specific inner product value corresponds to the largest possible angle between any two vectors in this space. Our implementation of the prototype pre-assignment slightly differs from (Yang et al., 2022) in two ways. First, we do not rely solely on the cross-entropy loss but instead incorporate the SupCon loss introduced earlier. Second, we do not assign the ETF prototypes randomly at the very beginning of training. Instead, we train the encoder with the SupCon for a few epochs and take advantage of the prototypes naturally formed by SupCon. Then, for each class, we assign the ETF prototype that is most closely aligned with its naturally formed prototype, and train the encoder to minimize the distance between the learned and pre-assigned prototype as:
$\begin{split}\mathcal{L}_{\text{{ETF}}}=\frac{1}{C^{0}}\sum_{i=1}^{C^{0}}(P_{c}-w_{c})^{2},\end{split}$ (2)
where, $P_{c}$ and $w_{c}$ are the assigned and learned prototypes of class $c$, and $C^{0}$ is the total number of classes in the base session.
@@
Including Pseudo-classes. Previous studies such as FACT (Zhou et al., 2022a) and SAVC (Song et al., 2023) introduced the concept of integrating pseudo-classes during the base session to serve as placeholders in the embedding space for novel classes. In SAVC, the pseudo-classes were generated by a pre-defined transformation, which was considered a more fine-grained variant of the original classes. In our work, we take a similar approach of including pseudo-classes from (Lee et al., 2020), which employs hard augmentations to transform the semantics of the sample and consider it as a pseudo-class. Let, $\mathcal{F}$ be a set of pre-defined (hard) augmentations for pseudo-class formation, and $x_{c,i}$ be a sample of class $c$. With the pseudo-class trick, we consider the transformation of the image $\mathcal{F}(x_{c,i})$ as an instance of a pseudo-class $(C^{0}*M+c)$, where $C^{0}$ is the total number of classes in base session, and $M$ is the pseudo-class multiplication factor. Such pseudo-classes can be seen as a fine-grained class derived from the original class. In our work, we use $M=2$, which doubles the total number of classes (including pseudo-classes) during base training. This trick does not include any new loss function.